Physics

An athlete starts at point and runs at a constant speed of 5.40 around a round track 250 in diameter

Find the component of this runner's average velocity between points A and B.

Find the component of this runner's average velocity between points A and B.

Find the component of this runner's average acceleration between points A and B.

Find the component of this runner's average acceleration between points A and B.
Answer #1
No point B is specified.
If B was the whole way around the track his displacement hasn't altered so his average velocity is zero.
(s2-s1)/t = (0)/t = 0
His velocity hasn't altered so the acceleration is zero.
(V2-v1)/t = (5.4-5.4)/t = 0/t = 0

If B was half way around the circle then the change in DISPLACEMENT is the diameter of the track
= 250m

And the change in velocity would be -5.4 - (5.4) = -10.8 m/s

All you need to do is work out the time from which to calculate velocity and acceleration
speed = distance / time
time = distance / speed (note, distance. NOT displacement)
= (Pi()*250 /2) / 5.4 = 72.68 s

So average velocity is displacement / time
= 250/72.68 = 3.44 m/s

Acceleration is change in velocity / change in time = -10.8/72.68
= - 0.15 m/s^2
answered by: John